OpenStudy (anonymous):

8(x + 1) - 3(x + 4) = 7(2 - x) x = -1 1/2 x = -1 1/4 x = 1 1/4 x = 1 1/2

OpenStudy (anonymous):

8x+8-3x-12=14-7x 5x-4=14-7x 5x+7x=14+4 12x=18 x=18/12 x=3/2 x= 1(1/2)

OpenStudy (anonymous):

x = 1 1/2 8x + 8 - 3x + 12 = 14 - 7x 8x - 3x - 8 + 12 5x + 4 5x+4 = 14 - 7x 5x = 10 - 7x 5x + 7x = 10 12x = 10 12/10 = x x= 6/5 6/5 = 1 and 1/2

OpenStudy (anonymous):

@MattNow14 u did a mistake in frst equation...

OpenStudy (anonymous):

@vedic what did I miss?

OpenStudy (anonymous):

@vedic , I see it now

OpenStudy (anonymous):

:):)